19. Randy walked at the rate of 5.2 km/h in a straight path from his campsite to a ranch. He returned immediately on horseback at the rate of 7.8 km/h Upon his returm, he found that he had been gone no more than 3.5 h. At most how far is it from his campsite to the ranch?
Answer #1

Let the ranch that Randy visited be x km from his campsite. He walked at a rate of 5.2km/h so that he would have walked x km in (x/5.2) hours. Further, Randy returned immediately on horseback at 7.8 km/h so that the x km ride would take (x/7.8 ) hours.

Now, since the journey from the campsite to the ranch and back took 3.5 hours, hence (x/5.2)+ (x/7.8 ) = 3.5 or, (5x/15.6 ) = 3.5 or, 5x = 3.5 * 15.6 = 54.6. Hence x = 54.6/5 = 10.92 km.

Thus, the ranch that Randy visited is 10.92 km from his campsite.